# Strassen $2\times 2$ Matrix Multiplication from a 3-dimensional Volume Form

### v1

@@ 5 Strassen algorithms
$\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})\\ +\frac{1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma(123)(i)}(a_{i}(v_{\sigma(i)})).$
@@
$\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})-\mathrm{tr}(a_{1}a_{2}a_{3})=\\ \frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)})(a_{i})$
@@
$c_{1,2}=v_{1}\lambda_{2}=\left(\begin{matrix}1&0\\ 0&0\end{matrix}\right),$ $c_{1,3}=v_{1}\lambda_{3}=\left(\begin{matrix}1&-1\\ 0&0\end{matrix}\right),$
$c_{2,3}=v_{2}\lambda_{3}=\left(\begin{matrix}0&0\\ 1&-1\end{matrix}\right),$ $c_{2,1}=v_{2}\lambda_{1}=\left(\begin{matrix}0&0\\ 0&1\end{matrix}\right),$
$c_{3,1}=v_{3}\lambda_{1}=\left(\begin{matrix}0&1\\ 0&1\end{matrix}\right),$ $c_{3,2}=v_{3}\lambda_{2}=\left(\begin{matrix}1&0\\ 1&0\end{matrix}\right).$
@@
These bilinear forms in the $a^{i,j}$ and $b^{i,j}$ are exactly the terms I, II, III, IV, V, VI, VII introduced in the original Strassen article [1]:
@@
exactly as originally stated by Strassen [1]. ∎

### v2

@@ 5 Strassen algorithms
$\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})\\ +\frac{1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma\cdot(1\,2\,3)(i)}(a_{i}(v_{\sigma(i)})).$
@@
$\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})-\mathrm{tr}(a_{1}a_{2}a_{3})=\\ \frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma\cdot(1\,2\,3)(i)})(a_{i})$
@@
$c_{1,2}=v_{1}\lambda_{2}=\biggl(\begin{matrix}1&0\\ 0&0\end{matrix}\biggr),$ $c_{1,3}=v_{1}\lambda_{3}=\biggl(\begin{matrix}1&-1\\ 0&\phantom{-}0\end{matrix}\biggr),$
$c_{2,3}=v_{2}\lambda_{3}=\biggl(\begin{matrix}0&\phantom{-}0\\ 1&-1\end{matrix}\biggr),$ $c_{2,1}=v_{2}\lambda_{1}=\biggl(\begin{matrix}0&0\\ 0&1\end{matrix}\biggr),$
$c_{3,1}=v_{3}\lambda_{1}=\biggl(\begin{matrix}0&1\\ 0&1\end{matrix}\biggr),$ $c_{3,2}=v_{3}\lambda_{2}=\biggl(\begin{matrix}1&0\\ 1&0\end{matrix}\biggr).$
@@
These bilinear forms in the $a^{i,j}$ and $b^{i,j}$ are exactly the terms I, II, III, IV, V, VI, VII introduced in the original Strassen article [Strassen]:
@@
exactly as originally stated by Strassen [Strassen]. ∎